### v2

# Independence and Induction in Reverse Mathematics

Abstract.
We continue the project of the study of reverse mathematics principles inspired by cardinal invariants. In this article in particular we focus on principles encapsulating the existence of large families of objects that are in some sense mutually independent. More precisely, we study the principle $\mathsf{MAD}$ stating that a maximal family of pairwise almost disjoint sets exists; and the principle $\mathsf{MED}$ expressing the existence of a maximal family of functions that are pairwise eventually different. We investigate characterisations of and relations between these principles and some of their variants. It will turn out that induction strength is an essential parameter in this context.
The investigators acknowledge the following partial support: F. Stephan’s reserach was supported by Singapore Ministry of Education AcRF Tier 2 grant MOE-000538-00 and AcRF Tier 1 grants A-0008454-00-00 and A-0008494-00-00, the grant A-0008494-00-00 also supported the research visits of R. Hölzl to the National University of Singapore and the past employment of D. Belanger at the National University of Singapore. C. T. Chong’s research was partially supported by NUS grant WBS E-146-00-0001-01.
1. Introduction
Reverse mathematics is the program of classifying the strength of mathematical statements relative to each other over a base system—see Simpson [6]. These mathematical statements are often referred to as “principles,” as in the Ramsey-style “combinatorial principles” that have captured such attention since the mid-1990’s.
Hölzl, Raghavan, Stephan and Zhang [2] introduced a number of principles modelled after notions of cardinal invariants from set theory. In this article we continue and extend this study with particular focus on principles that are in some sense related to the mutual independence of objects. More precisely, we use the notion of weakly represented families of objects introduced in the above article [2] to express the existence of large families of objects and then investigate from a reverse-mathematical point of view the strength of principles that assert the existence of maximal families of objects that are “independent” of each other in one sense or another. We begin by recalling some basic facts in reverse mathematics.
2. Preliminaries
We are interested in models in the language of second-order arithmetic. These take the form $\mathcal{M}=(M,\mathcal{S},+,\cdot,0,1,\in)$ where $M$ is the first-order part and $\mathcal{S}$ is a subset of the power set of $M$. Typically we suppress the operations and write simply $\mathcal{M}=(M,\mathcal{S})$. We call the elements of $M$ numbers, with the idea that they stand in for the natural numbers $\omega$ and obey many of the same axioms. In this article we resist the convention of writing “sets” to mean the elements of $\mathcal{S}$, because we will often need to discuss sets $A\subseteq M$ which are possibly outside of $\mathcal{S}$. Call a formula a $\Sigma^{A}_{n}$ formula if it is $\Sigma_{n}$ while using the given set $A$ as a second-order parameter (which we also call an oracle). Call a set $B\subseteq M$ a $\Sigma^{A}_{n}$ set if $B$ is definable by a $\Sigma_{n}^{A}$ formula. Formulas with multiple oracles are allowed; call a formula or set $\Sigma^{0}_{n}$ if it is $\Sigma^{A_{1},\ldots,A_{k}}_{n}$ for some $A_{1},\ldots,A_{k}\in\mathcal{S}$; this includes the case where $k=0$ and the list is empty. Make all the analogous definitions with $\Pi$ and $\Delta$ in place of $\Sigma$.
Here are some of the axioms we will consider. Let $P^{-}$ be the axioms of Peano arithmetic not including the induction scheme. Let $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ denote the induction scheme for $\Sigma^{0}_{n}$ formulas, and let $\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ denote the $\Sigma^{0}_{n}$-bounding scheme. Kirby and Paris [3] proved over the system $P^{-}$ that $\mathsf{B}\mathrm{\Sigma}_{n+1}^{0}\rightarrow\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ and $\mathsf{I}\mathrm{\Sigma}_{n}^{0}\rightarrow\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ for $n\geq 1$, and that these implications are strict. (In their account there were no second-order parameters, but the proof is the same.) The system $\mathsf{RCA}_{0}$ consists of $P^{-}$ together with $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ and the comprehension scheme for $\Delta^{0}_{1}$ sets. In particular, if $\mathcal{M}=(M,\mathcal{S})$ is a model of $\mathsf{RCA}_{0}$, then $S$ is closed under Turing reducibility.
[…]
In this paper, we consider principles related to set theory’s cardinal invariants, in the context of $\mathsf{RCA}_{0}$. In particular, we are interested in the role played by the inductive strength of a model of $\mathsf{RCA}_{0}$ in relation to these principles. This naturally requires an analysis of nonstandard models.
The families of functions and families of sets that we consider are the same as those considered by Hölzl, Raghavan, Stephan and Zhang [2], and informally speaking consist of those functions that can be written as the “rows” $\Psi_{e}$ of some set $\Psi$, with the following details: we require that $\Psi$ be $\Sigma^{0}_{1}$ (with parameters), as opposed to being an element of $\mathcal{S}$; we require that each row $\Psi_{e}$ be well-defined as a function, but not necessarily total; and all rows $\Psi_{e}$ which are not total are simply ignored and do not contribute to the family. More formally, we will work with the following definition.
Definition 1 (Weakly represented families).
Let $\Psi\subseteq M^{2}\rightarrow M$ be a $\Sigma^{0}_{1}$ (with set parameters) partial function, and for each $e$ define $\Psi_{e}=\{\langle x,y\rangle\colon\Psi(e,x)=y\}$. View $\Psi_{e}$ as a partial function $M\rightarrow M$. Define =${\mathcal{F}}=\{\Psi_{e}\colon\Psi_{e}\leavevmode\nobreak\ \textnormal{is\leavevmode\nobreak\ total}\}$. Then ${\mathcal{F}}$ is a family of total functions $M\rightarrow M$; we say that ${\mathcal{F}}$ is weakly represented by $\Psi$. In this case we call ${\mathcal{F}}$ a weakly represented family.
By identifying a set $A$ with its characteristic function, we can naturally speak of a family ${\mathcal{F}}$ of sets $A\subseteq M$ being weakly represented. If ${\mathcal{F}}$ is a weakly represented family (of functions or of sets), then each element $f\in{\mathcal{F}}$ is $\Delta^{0}_{1}$ relative to sets in $S$ and therefore is in $\mathcal{S}$. By considering the model $\mathcal{M}$’s version of the universal oracle Turing machine, we obtain the following result.
Lemma 2.
For each $A\in\mathcal{S}$ there exists a $\Sigma^{A}_{1}$ partial function $\Phi^{A}\colon M^{2}\rightarrow M$ which is universal in the sense that for every $\Sigma^{A}_{1}$ partial function $\Psi^{A}\colon M^{2}\rightarrow M$ there is a $\Delta_{1}^{A}$ function $f$ satisfying $\Psi^{A}_{e}=\Phi^{A}_{f(e)}$.
We use ${\mathcal{F}}^{A}_{\mathrm{Tot}}$ to denote the family of functions weakly represented by $\Phi^{A}$. Then every weakly represented family ${\mathcal{F}}$ is contained in ${\mathcal{F}}^{A}_{\mathrm{Tot}}$ for some $A\in\mathcal{S}$. Lastly, we fix a definition of finite that is suitable for weakly represented families.
Definition 3.
A weakly represented family ${\mathcal{F}}$ is called finite if there is a $\Psi$ which weakly represents it, and an $\mathcal{M}$-finite set $E$ such that =${\mathcal{F}}=\{\Psi_{e}\colon e\in E\}$. Otherwise, ${\mathcal{F}}$ is called non-finite.
In the absence of stronger induction axioms, a family ${\mathcal{F}}$ which is non-finite in this sense does not satisfy certain intuitive properties of infinity. As an example: If $\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ holds, then the family ${\mathcal{F}}$ conisting of functions $f_{e}(x)=e$ if $\Phi_{e}^{A}(x){\downarrow}$ for $e<e_{0}$ is bounded, but in general not finite; this is because $\{e<e_{0}\colon\Phi_{e}^{A}\textnormal{is\leavevmode\nobreak\ total}\}$ is in general not $M$-finite. We bring this up to explain the presence of the unusual term non-finite, as well as to point out that certain of our results below depend on this definition, and could end up different if some other notion of finite were used.
The rest of this paper is structured as follows. Section 3 deals with the principle $\mathsf{DOM}$ which expresses essentially that there is a function dominating every total function in a weakly represented family; we prove a version of Martin’s theorem [5] characterizing such functions in terms of high Turing degrees. Section 4 is a technical section where we carry out a gener al $0^{\prime\prime}$ tree argument for use in later sections, to construct new weakly represented families. Section 5 deals with the principle $\mathsf{MAD}$, which says that there is a maximal almost disjoint (MAD) family of sets. We obtain results comparing $\mathsf{MAD}$ with $\mathsf{DOM}$, and with $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$. In Section 6, we investigate the principles $\mathsf{MED}$, which says there is a maximal eventually different (MED) family of functions, as well as $\mathsf{AVOID}$, which says essentially that there is a function eventually different (ED) from all total functions in a weakly represented family; we obtain results about the relationship between the two. Section 7 covers the principle $\mathsf{NDF}$, which says essentially that there is a hyperimmune Turing degree. Finally, Section 8 is concerned with the principle $\mathsf{BI}$, which says that there is a bi-immune Turing degree; we use the existence of a low bi-immune Turing degree to obtain a conservation result.
3. Dominating families of functions
Statement 4 (Domination principle, $\mathsf{DOM}$).
Given any weakly represented family $\mathcal{F}$ of functions, there exists a function $g$ such that $g$ dominates $\mathcal{F}$, that is, for every $f\in\mathcal{F}$, we have $g(x)>f(x)$ for all sufficiently large $x$.
In an $\omega$-model—meaning a model $(M,\mathcal{S})$ in which $M$ is the true natural numbers—if $X\in\mathcal{S}$ computes a function which dominates every total recursive function, then $X$ is generalized high, that is, $X^{\prime}\equiv_{T}(X\oplus\emptyset^{\prime})^{\prime}$. This was first proved by Martin [5]. The phenomenon is reflected in $\mathsf{RCA}_{0}$ by the following proposition, proved essentially in Chong, Qian, Slaman and Yang [1, Lemma 2]:
Proposition 5 (Martin’s Theorem, formalized).
Over $\mathsf{RCA}_{0}$, the following are equivalent:
  1. (i)
    $\mathsf{DOM}$
  2. (ii)
    $\mathsf{HIGH}{:}$ For every $A$, there is a $B$ such that every $\Sigma^{A}_{2}$ is $\Delta^{B}_{2}$.
  3. (iii)
    For each $A$, every $\Sigma^{A}_{2}$ set is $\Delta^{B}_{2}$ for some $B$.
Proof.
(i $\Rightarrow$ ii) Fix $A$ and let $f$ be a total function which dominates all total $\Delta_{1}^{A}$ functions. Take any $C$ which is $\Pi_{2}^{A}$, and let $\psi^{A}$ be the $\Delta_{0}^{A}$ formula satisfying
$x\in C\iff(\forall y)(\exists z)[\psi^{A}(x,y,z)].$
Define $g(x,y)$ to be the partial $\Sigma_{1}^{A}$ function mapping each pair $\langle x,y\rangle$ to the least $z$ satisfying $\psi^{A}(x,y,z)$. Then for a given $x$, we have $x\in C$ iff the function $\lambda y.g(x,y)$ is total $\Delta_{1}^{A}$. Hence,
$\begin{split}x\in C\iff(\exists y_{0})(\exists z_{0})\left[\begin{array}[]{c}(\forall y>y_{0})(\exists z<f(y))[\psi^{A}(x,y,z)]\\ \textnormal{and}\leavevmode\nobreak\ (\forall y\leq y_{0})(\exists z<z_{0})[\psi^{A}(x,y,z)]\end{array}\right],\end{split}$
meaning $C$ is $\Sigma_{2}^{f}$ and hence $\Delta_{2}^{f}$.
(ii $\Rightarrow$ iii) Immediate; the difference between (ii) and (iii) is that whereas in (ii) a single $B$ is assigned to each $A$, in (iii) $B$ is allowed to vary with each $\Sigma^{A}_{2}$-definable set.
(iii $\Rightarrow$ i) Suppose ${\mathcal{F}}$ is a weakly represented family of functions. Then ${\mathcal{F}}\subseteq{\mathcal{F}}^{A}_{\mathrm{Tot}}$ for some $A$; so it suffices to exhibit some $g$ which dominates all $f\in F^{A}_{\mathrm{Tot}}=\{\Phi_{e}^{A}\colon\Phi_{e}^{A}$ is total}}\}}.
Since the predicate “$\Phi_{e}^{A}$ is total” is $\Pi^{A}_{2}$, there exists by (iii) a $B$ and a $\Sigma^{B}_{0}$ formula $\psi$ satisfying
$\Phi^{A}_{e}\textnormal{\leavevmode\nobreak\ is\leavevmode\nobreak\ total}\iff(\exists s)(\forall t)[\psi^{B}(e,s,t)].$
Then for every input $x$ let $g(x)$ be the least $t_{0}>x$ such that
$(\forall e<x)\left[\textnormal{either}\leavevmode\nobreak\ \Phi_{e,t_{0}}^{A}(x){\downarrow}\leavevmode\nobreak\ \textnormal{or}\leavevmode\nobreak\ (\forall s<x)(\exists t<t_{0})[\neg\psi^{A}(e,s,t)]\right].$
Then $g$ is well-defined, $\Sigma_{1}^{B}$ and total. By definition it dominates each total $\Delta_{1}^{A}$ function. ∎
This formalized version of Martin’s Theorem allows for a short proof of the previously known fact that $\mathsf{DOM}$ and $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ together imply full arithmetical induction.
Theorem 6 (Hölzl, Raghavan, Stephan, Zhang [2]).
$\mathsf{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ for every $n\geq 1$.
We will use the following well-known fact in the new proof.
Lemma 7.
The following holds for every model $(M,\mathcal{S})$ of $\mathsf{RCA}_{0}$:
  1. (1)
    $\mathsf{I}\mathrm{\Sigma}_{n}^{0}$ iff every $\Sigma^{0}_{n}$ set is regular iff every $A^{(n)}$ is regular for $A\in\mathcal{S}$,
  2. (2)
    $\mathsf{B}\mathrm{\Sigma}_{n}^{0}$ iff every $\Delta^{0}_{n}$ set is regular iff every $B\leq_{T}A^{(n-1)}$ is regular for $A\in\mathcal{S}$.
Proof of Theorem 6.
Fix $(M,\mathcal{S})\models\text{RCA}_{0}+\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and let $A_{0}\in\mathcal{S}$. Use Proposition 5 to obtain a sequence $\{A_{n}:n\in\omega\}$ such that $A_{n+1}$ is high relative to $A_{n}$. In particular, the double jump $A_{n}^{\prime\prime}$ is $\Delta_{2}^{A_{n+1}}$ and therefore regular by $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ and Lemma 7. Since $A_{0}^{\prime\prime}\leq_{T}A_{1}^{\prime}$, we also have $A_{0}^{\prime\prime\prime}\leq_{T}A_{1}^{\prime\prime}$. Hence by transitivity ($\mathsf{B}\mathrm{\Sigma}_{1}$ relative to $A_{2}^{\prime}$), we have $A_{0}^{\prime\prime\prime}\leq_{T}A_{2}^{\prime}$ and hence that $A_{0}^{\prime\prime\prime}$ is regular. Iterating this argument for larger and larger $n$, we find that $A_{0}^{(n)}\leq_{T}A_{n-1}^{\prime}$ and $A_{0}^{(n)}$ is regular for all $n\in\omega$. The result follows from Lemma 7 and the fact that $A_{0}$ was arbitrary. ∎
Thus it is all the more surprising that, by the following theorem, $\mathsf{DOM}$ does not imply $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$. The following theorem also gives an alternative (but similar) proof that $\mathsf{HIGH}$ does not imply $\mathsf{ACA}_{0}$, by way of conservation in place of cone avoidance. Recall that a system $T_{1}$ is $\Pi^{1}_{1}$-conservative over a system $T_{2}$ if $T_{1}\supset T_{2}$ and every $\Pi^{1}_{1}$-sentence provable in $T_{1}$ is provable in $T_{2}$. By a well-known observation of Harrington, one can establish $\Pi^{1}_{1}$-conservation by showing that every countable model $\mathcal{M}=(M,\mathcal{S})$ of $T_{2}$ may be extended to a model $\mathcal{M}^{\prime}=(M,\mathcal{S}^{\prime})$ of $T_{1}$ having the same first-order part $M$ and such that $\mathcal{S}\subset\mathcal{S}^{\prime}$.
Theorem 8.
$\mathsf{RCA}_{0}+\mathsf{DOM}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}$.
Proof.
Let $\mathcal{M}_{0}=(M,\mathcal{S}_{0})$ be a countable model of $\mathsf{RCA}_{0}$ topped by a function $f_{0}\in\mathcal{S}$; that is, every $A\in\mathcal{S}$ is Turing-reducible to $f_{0}$. We will construct a sequence of functions $(f_{n})_{n<\omega}$, and with it a sequence of countable models $(\mathcal{M}_{n})_{n<\omega}$ given by
$\mathcal{M}_{n+1}=\mathcal{M}_{n}[f_{n+1}]=(M,\mathcal{S}_{n+1}),$
where $\mathcal{S}_{n+1}$ is the closure of $\mathcal{S}\cup\{f_{n+1}\}$ under $\Delta^{0}_{1}$ comprehension. We ensure that each $f_{n+1}$ dominates ${\mathcal{F}}_{\mathrm{Tot}}^{f_{n}}$ and each $\mathcal{M}_{n}$ satisfies $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ (and, in particular, $\mathcal{M}_{n}$ is topped by $f_{n}$). Then the limiting model $\mathcal{M}_{\omega}=(M,\bigcup_{n<\omega}\mathcal{S}_{n})$ will satisfy $\mathsf{RCA}_{0}+\mathsf{DOM}$, proving the theorem by Harrington’s method.
For the inductive step, begin with the countable model $\mathcal{M}_{n}\models\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ topped by $f_{n}$, and consider the class ${\mathcal{F}}_{\mathrm{Tot}}^{f_{n}}$. Since $\mathcal{M}_{n}$ is countable, we may (externally to the model) let $(g_{i})_{i<\omega}$ be a list of all $\Delta_{1}^{f_{n}}$ total functions; let $(\varphi_{i}^{f_{n}})_{i<\omega}$ be a list of all $\Sigma_{1}^{f_{n}}$ formulae; and let $(c_{k})_{k<\omega}$ be an increasing sequence that is cofinal in $M$. We construct $f_{n+1}$ by initial segments $\sigma_{k}\in M^{<M}$. Beginning with $\sigma_{0}$ as the empty string, for each $k\in\omega$ we choose the next $\sigma_{k+1}$ to satisfy
  • (i)
    $\sigma_{k+1}$ is a proper extension of $\sigma_{k}$,
  • (ii)
    $|\sigma_{k+1}|>c_{k}$,
  • (iii)
    $\sigma_{k+1}(x)\geq g_{i}(x)$ whenever $i<k$ and $|\sigma_{k}|<x\leq|\sigma_{k+1}|$, and
  • (iv)
    either $\mathcal{M}_{n}\models\varphi_{k}^{f_{n}}(\sigma_{k+1})$ or $\mathcal{M}_{n}\models(\forall\tau\supseteq\sigma_{k+1})[\neg\varphi_{k}^{f_{n}}(\tau)]$.
Because $k$ is always in $\omega$, some suitable $\sigma_{k+1}$ exists at every stage. This completes the construction.
The requirements (i) and (ii) together ensure that $f$ is a well-defined, total function (rather than having a proper cut as its domain). Item (iii) ensures that $f_{n+1}$ dominates each function in $\mathcal{M}_{n}$, and (iv) is a form of “forcing the jump” that ensures that $\mathcal{M}_{n}[f_{n+1}]$ models $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$. This completes the proof. ∎
4. Handling index sets
This technical section lists some methods for producing new weakly represented families ${\mathcal{G}}$ from known families ${\mathcal{F}}$, in models of $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$. These take the form of sufficient conditions on the set $B$ in the definition ${\mathcal{G}}=\{\Psi_{e}\colon e\in B\}$, where ${\mathcal{F}}$ is weakly represented by $\Psi$. However, the constructions are not as simple as just restricting $\Psi$ to the rows in $B$, because we would like $B$ to be of higher complexity than that would allow. Instead, we use a conventional $0^{\prime\prime}$ infinite injury argument—reminiscent of, for instance, Yates [7]—to get a new $\Theta$ which weakly represents ${\mathcal{G}}$, and which can be viewed as a scrambled version of $\Psi$ restricted to $B$.
We begin with an easy but helpful lemma; here, we let
$\mathrm{rt}B=\{\sigma\in 2^{<M}\colon\sigma(b)=0\rightarrow b\in B\},$
that is, $\mathrm{rt}B$ consists of the binary strings which are either substrings of the complement $\overline{B}$’s characteristic function, or lexicographically to the right of that characteristic function.
Lemma 9.
Let $(M,\mathcal{S})$ be a model of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ and assume that $B\subseteq M$ is $\Pi^{A}_{2}$ for some $A$. Then $\mathrm{rt}B$ is $\Pi^{A}_{2}$.
Proof.
Fix $A$ such that $B$ is $\Pi_{2}^{A}$. Then $B$ is $\Pi_{1}^{A^{\prime}}$. Then clearly $\mathrm{rt}B$ is $\Pi_{1}^{A^{\prime}}$. So $\mathrm{rt}B$ is $\Pi_{2}^{A}$. ∎
Next we present three new methods of representing families of functions under different assumptions; in each case, the $\Theta$ we are constructing can be thought of as a restricted and scrambled version of the input representation $\Psi$.
Lemma 10 ($\Pi^{0}_{2}$ weak representation).
Suppose $\mathcal{M}=(M,\mathcal{S})$ is a model of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ and $\Psi$ is $\Sigma^{0}_{1}$. If $B\subseteq M$ is $\Pi^{0}_{2}$, then the family $\{\Psi_{e}\colon e\in B\textnormal{ and }\Psi_{e}\textnormal{ is total}\}$ is weakly represented.
Proof.
Let $\mathrm{rt}B$ be as in Lemma 9. Since $\mathrm{rt}B$ is $\Pi^{0}_{2}$, it has a $\Delta^{0}_{1}$ approximation $(\sigma_{s})_{s\in M}$ satisfying $\sigma\in\mathrm{rt}B$ iff $(\exists^{\infty}s)[\sigma=\sigma_{s}]$. Begin the construction of a new $\Sigma^{0}_{1}$ set $\Theta$ by assigning to each $\sigma\in 2^{<M}$ a row $\Theta_{\sigma}$, while leaving $\mathcal{M}$-infinitely many rows of $\Theta$ unassigned. Then proceed by stages as follows:
At even stages $2s$, for each $\tau$ strictly to the right of $\sigma_{s}$, abandon its currently assigned row $\Theta_{\tau}$ and assign to it a new, fresh row (which now takes the name $\Theta_{\tau}$). Leave $\mathcal{M}$-infinitely many rows of $\Theta$ still unassigned.
At odd stage $12s+$ 1, for each $e<|\sigma_{s}|$ for which $\sigma_{s}(e)=0$, make $\Theta_{\sigma_{s}\upharpoonright e}$ imitate $\Psi_{e}$ for $s$-many steps of computation.
This completes the construction. Clearly $\Theta$ is a partial function, and every total row of $\Theta$ is the row $\Theta_{\sigma}$ eventually assigned to some $\sigma$ which is an initial segment of the complement $\overline{B}$’s characteristic function. And on the other hand, such a $\Theta_{\sigma}$ is total if and only if $\sigma(e)=0$ (meaning $|\sigma|-1\in B$) and $\Psi_{|\sigma|-1}$ is total. Hence $\Theta$ weakly represents the desired family of functions. ∎
Lemma 11 ($\Sigma^{0}_{2}$ weak representation).
As in Lemma 10, but with $B$ being $\Sigma^{0}_{2}$ instead of $\Pi^{0}_{2}$.
This can be seen by almost the same proof as above, but using $\overline{B}$ as the $\Pi^{0}_{2}$ set and accordingly replacing “$\sigma_{s}(e)=0$” by “$\sigma_{s}(e)=1$” in the odd stages.
Lemma 12 ($\Pi^{0}_{2}$ dependent weak representation).
Suppose $\mathcal{M}=(M,\mathcal{S})$ models ${\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}}$ and that $\varphi$ is a $\Pi^{0}_{2}$ predicate. Then
$\{\Psi_{e_{i}}\colon e_{i}\leavevmode\nobreak\ is\leavevmode\nobreak\ least\leavevmode\nobreak\ such\leavevmode\nobreak\ that\leavevmode\nobreak\ \varphi(\langle e_{0},\ldots,e_{i-1}\rangle)\textnormal{ holds}\}$
is a weakly represented family.
In other words: $e_{0}$ is least such that $\varphi(\langle e_{0}\rangle)$ holds, then $e_{1}$ is least such that $\varphi(\langle e_{0},e_{1}\rangle$) holds, and so on. It is possible that $e_{i}$ is defined for all $i\in M$; it is possible that there is no $i$ at all, or a maximum $i$ for which $e_{i}$ is defined; and it is also possible that the set of $i$ forms a proper cut.
Proof.
Associate to each $\sigma\in 2^{<M}$ the sequence $\langle e_{0}^{\sigma},\ldots,e_{k}^{\sigma}\rangle\in M^{<M}$ where
  • $e_{0}^{\sigma}$ is the number of $1$’s occurring before the first 00 in $\sigma$, and
  • $e_{i+1}^{\sigma}$ is the number of $1$’s occurring between the $(i+1)$-th and the $(i+2)$-th 00 in $\sigma$; and
  • for each $\sigma$, we use $k$ to mean the largest number for which $e_{k}^{\sigma}$ is defined.
Then the set of $\sigma$ for which $\mathcal{M}\models\varphi(\langle e_{0}^{\sigma},\ldots,e_{k}^{\sigma}\rangle)$ holds is closed under initial segment — in other words, it forms a binary tree—it is $\Pi^{0}_{2}$ definable, and its leftmost infinite path $\sigma$ corresponds to the sequence $\langle e_{0},e_{1},\ldots\rangle$ described in the lemma’s statement. By applying the argument of Lemma 10 to this tree and its $e_{i}$, we obtain the required $\Theta$. ∎
5. Maximal, almost disjoint families of sets
Two sets $A,B\subseteq M$ are said to be almost disjoint if $A\cap B$ is $\mathcal{M}$-finite. A family ${\mathcal{F}}$ of subsets of $M$ is almost disjoint if any two distinct elements of ${\mathcal{F}}$ are almost disjoint. ${\mathcal{F}}$ is called maximal almost disjoint if it is non-finite, almost disjoint and not properly contained in any almost disjoint family.
Statement 13 (Maximal Almost Disjoint Family principle, $\mathsf{MAD}$).
There is a non-finite, weakly represented family of infinite sets that is maximal almost disjoint.
Hölzl, Raghavan, Stephan and Zhang showed in [2] that $\mathsf{MAD}$ is equivalent to $\neg\mathsf{DOM}$ in $\omega$-models. We extend this result to all models of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
Theorem 14.
$\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}\vdash\mathsf{DOM}\leftrightarrow\neg\mathsf{MAD}$.
Proof.
($\mathsf{DOM}\rightarrow\neg\mathsf{MAD}$) Fix a model $\mathcal{M}=(M,\mathcal{S})\models\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{DOM}$, and any non-finite almost disjoint family ${\mathcal{F}}$ which is weakly represented, say by $\Psi^{A}$ for some $A\in\mathcal{S}$ (note that since $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ holds in the model, ${\mathcal{F}}$ is in fact $\mathcal{M}$-infinite). The set
[…]
is $\Pi^{A}_{2}$, and hence by $\mathsf{DOM}$ and Proposition 5 it is $\Delta^{f}_{2}$ for some $f\in\mathcal{S}$. Let $(E_{s})_{s\in M}$ be a $\Delta^{f}_{1}$ approximation to $E$.
Define a new set $B=\{b_{0}<b_{1}<\ldots\}$ as follows. Let $b_{0}$ be any number. To define $b_{k+1}$, search for a pair $x,s>b_{k}$ satisfying
$(\forall e<k)[\textnormal{either}\leavevmode\nobreak\ \Psi_{e,s}^{A}(x){\downarrow}=0\leavevmode\nobreak\ \textnormal{or}\leavevmode\nobreak\ e\not\in E_{s}],$
and let $b_{k+1}$ be this $x$. Since $B$ is being enumerated in increasing order and in a $\Sigma^{f}_{1}$ way, we know by $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ that $B\in\mathcal{S}$ and (using the definition of $E$) that $b_{k}$ is defined for every $k\in M$. And since $(E_{s})_{s\in M}$ reaches a pointwise limit, $\mathcal{M}\models(\forall e)(\forall^{\infty}k)[b_{k}\not\in\Psi_{e}^{A}]$. Hence, if we let $\Theta$ equal $\Psi$ prepended with a row for $B$’s characteristic function, we see that ${\mathcal{F}}\cup\{B\}$ is a weakly represented family strictly larger than ${\mathcal{F}}$. Now fix any $C\in{\mathcal{F}}$; we claim that $B$ and $C$ are almost disjoint. Suppose for a contradiction that $B\cap C$ were $\mathcal{M}$-infinite, and let $\Psi^{A}_{e}$ be $C$’s characteristic function. Then from the definition of $b_{k}$ we know that $e\not\in E_{s}$ for unboundedly many $s$, and so in the limit we have $e\not\in E$. But this contradicts the definition of $E$. Therefore ${\mathcal{F}}$ is contained in the strictly larger almost disjoint family ${\mathcal{F}}\cup\{B\}$, and in particular ${\mathcal{F}}$ is not a MAD family. Since ${\mathcal{F}}$ was arbitrary, we conclude $\neg\mathsf{MAD}$ holds.
($\neg\mathsf{MAD}\rightarrow\mathsf{DOM}$) Fix any model $\mathcal{M}=(M,\mathcal{S})\models\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{MAD}$, fix any $A\in\mathcal{S}$, and let $\Phi^{A}$ be universal as in Lemma 2. Consider the formula $\psi$ which takes a tuple $\langle e_{0},\ldots,e_{k}\rangle$ of indices as input and expresses that
  • (a)
    each $\Phi^{A}_{e_{i}}$ is the characteristic function of an $\mathcal{M}$-infinite set—call it $C_{i}$;
  • (b)
    these $C_{i}$’s are pairwise disjoint;
  • (c)
    the union $D_{k}=\bigcup_{i<k}C_{i}$ is $\mathcal{M}$-coinfinite; and
  • (d)
    for all $x$ and $e<i<k$, either $\Phi^{A}_{e}$ is not total or $\Phi^{A}_{e}(x)$ is less than the $x$-th smallest element of $\overline{D_{i}}$.
These statements are $\Pi^{0}_{2}$; for (a)–(c) that is straightforward; for (d) consider the formula
$(\forall x)(\forall i<k)(\forall s)(\exists t>s)(\exists y)(\forall z<y)(\forall e<i)[z\in D_{i,t}\wedge(\Phi^{A}_{e,t}(x){\downarrow}<y\vee e\in\mathrm{Tot}^{A}_{t})],$
where $(\mathrm{Tot}^{A}_{t})_{t}$ is the usual $A$-computable approximation to $\mathrm{Tot}^{A}=\{e\in M\colon\Phi_{e}^{A}$ is total}}\}}.
Thus, applying Lemma 12, we obtain some $\Theta$ weakly representing
=${\mathcal{F}}=\{\Phi^{A}_{e_{i}}\colon e_{i}\in\langle e_{0},e_{1},\ldots\}$
, where $\langle e_{0},e_{1},\ldots\rangle$ is lexicographically least satisfying both $\psi$ and the relation $e_{0}<e_{1}<\cdots$.
Now ${\mathcal{F}}$ forms a partition of $M$ into $\mathcal{M}$-infinite sets, and hence is an almost disjoint family. Since $\mathcal{M}\models\neg\mathsf{MAD}$, there is a non-finite $B\in\mathcal{S}$ almost disjoint from each $C_{i}$. By $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ this $B$ is almost disjoint from each $D_{i}$, meaning all but $\mathcal{M}$-finitely much of $B$ is contained in each complement $\overline{D_{i}}$. Hence the function $p_{B}$ which maps $x$ to the $x$-th element of $B$ dominates all total $\Phi_{e}^{A}$ for $e<i$. By allowing $i$ to vary, we see that $p_{B}$ dominates all total $\Delta^{A}_{1}$ functions. Since $A\in\mathcal{S}$ was arbitrary, we conclude that $\mathsf{DOM}$ holds. ∎
An alternate proof of $\neg\mathsf{MAD}\rightarrow\mathsf{DOM}$ might construct the partition directly using a $0^{\prime\prime}$ tree argument similar to the proof of Lemma 10. A natural next question is whether $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is necessary for the equivalence $\neg\mathsf{MAD}\leftrightarrow\mathsf{DOM}$ to hold. The answer is yes, although this is in part an artefact of our choice of definition for finite weakly represented families.
Proposition 15.
$\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}\vdash\mathsf{MAD}$.
Proof.
Fix a model $\mathcal{M}=(M,\mathcal{S})$ and let $A\in\mathcal{S}$ be such that there is a $\Pi^{A}_{2}$ set $B$ which is bounded from above but not $\mathcal{M}$-finite; such $A,B$ are guaranteed to exist by $\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ and Lemma 7. It is routine to use the $\Pi^{A}_{2}$ formula defining $B$ and its upper bound $a\in M$ to define a partition $C_{0},\ldots,C_{a-1}\in\mathcal{S}$ of $M$ such that $C_{i}$ is $\mathcal{M}$-infinite iff $i\in B$. Let $\Psi^{A}\colon M^{2}\rightarrow M$ be the $\Sigma^{A}_{1}$ partial function
$\Psi^{A}(i,s)=\begin{cases}1&\textnormal{if }i<a\textnormal{ and }s\in C_{i},\\ 0&\textnormal{if }i<a\textnormal{ and }s\not\in C_{i}\textnormal{ and }(\exists s>t)[t\in C_{i}],\\ {\uparrow}&\textnormal{otherwise.}\\ \end{cases}$
Then $\Psi^{A}$ weakly represents the family =${\mathcal{F}}=\{C_{i}\colon i\in B\}$. Since the $C_{i}$ form a partition, ${\mathcal{F}}$ is almost disjoint; we claim that ${\mathcal{F}}$ is MAD. To see this, suppose for a contradiction that some $D\in\mathcal{S}$ is $\mathcal{M}$-infinite, but $D\cap C_{i}$ is $\mathcal{M}$-finite for every $i\in B$. Then $D\cap C_{i}$ is $\mathcal{M}$-finite for every $i<a$. That is, we have $(\forall i<a)(\exists s)(\forall t>s)[t\not\in C_{i}\cap D]$. Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the two leading quantifiers yields $(\exists s)(\forall i<a)(\forall t>s)[t\not\in C_{i}\cap D]$. Since the $C_{i}$ form a partition, this implies that $D$ is $\mathcal{M}$-finite, a contradiction.
To complete the argument, we claim that ${\mathcal{F}}$ is a non-finite family. Assume for a contradiction that it is finite, that is, that there exists some other $\Theta$ weakly representing ${\mathcal{F}}$ such that there is an $\mathcal{M}$-finite set $E$ with =${\mathcal{F}}=\{\Theta_{e}\colon e\in E\}$. Then, for each $e\in E$, we can find one $w_{e}\in\Theta_{e}$ and using $A$ as an oracle the unique $i$ such that $w_{e}\in C_{i}$. But this gives an effective procedure for determining which $i<a$ are elements of $B$. Thus $B$ would be $\mathcal{M}$-finite, contradicting our assumption. ∎
6. Avoidance and eventual difference
Two functions are eventually different if there is an upper bound on the inputs where they are equal; they are eventually equal if there is an upper bound on the inputs where they are different.
Statement 16 (Avoidance principle, $\mathsf{AVOID}$).
Given any weakly represented family of functions $\mathcal{F}$, there exists a function $g$ that is eventually different from all $f\in\mathcal{F}$.
Statement 17 (Maximal Eventually Different Family principle, $\mathsf{MED}$).
There is a weakly represented family ${\mathcal{F}}$ of functions such that
  • any two functions in ${\mathcal{F}}$ are either equal or eventually different, and
  • every function coincides infinitely often with some function in ${\mathcal{F}}$.
Theorem 18.
$\mathsf{RCA}_{0}\vdash\mathsf{MED}\rightarrow\neg\mathsf{AVOID}$.
Proof.
Clearly, any family witnessing $\mathsf{MED}$ is a counterexample to $\mathsf{AVOID}$. ∎
The following is a reformulation of a previously known result in the terminology of reverse mathematics; we include our own proof since as far as we know it is more direct than others that have appeared in print.
Proposition 19 (Kjos-Hanssen, Merkle, Stefan [4]).
$\mathsf{RCA}_{0}\vdash(\mathsf{DOM}\vee\mathsf{DNR})\leftrightarrow\mathsf{AVOID}$.
Proof.
That $\mathsf{DOM}$ implies $\mathsf{AVOID}$ is immediate from the definitions. To see that $\mathsf{DNR}$ implies $\mathsf{AVOID}$, first suppose that $f$ is a DNR function relative to a given $A\in\mathcal{S}$, that is, that $(\forall e)[\Phi^{A}_{e}(e)\neq f(e)]$. Let $\Phi^{A}$ be universal in the sense of Lemma 2, and for each $e$ let $p_{e}$ be the $e$-th prime number, starting with $p_{0}=2$ and possibly continuing into the nonstandard numbers. Define a function $g\leq_{T}f$ as follows. Given $e$, find a sequence ${i_{0},\ldots,i_{e}}$ of indices of partial functions, for each $j$ satisfying
$\Phi^{A}_{i_{j}}(i_{j})=\begin{cases}0&\textnormal{if }\Phi^{A}_{j}(e)=0,\\ k&\textnormal{if }k\textnormal{ is greatest such that }p_{j}^{k}\textnormal{ divides }\Phi^{A}_{j}(e),\\ \end{cases}$
and let $g(e)=p_{0}^{f(i_{0})}p_{1}^{f(i_{1})}\cdots p_{e}^{f(i_{e})}$. This ensures that $g(e)\neq\Phi^{A}_{j}(e)$ for every ${j\leq e}$; and thus for every $j\in M$, we have that $g$ is eventually different from $\Phi^{A}_{j}$.
[…]
$h(e)=\begin{cases}\Phi_{e,g(e)}&\textnormal{if }\Phi_{e,g(e)}{\downarrow},\\ 0&\textnormal{otherwise,}\\ \end{cases}$
is total and $\mathcal{M}$-infinitely equals $f$. This contradiction concludes the proof. ∎
To conclude this section, we will consider the role of $\mathsf{MED}$ in models of $\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$. In the presence of $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$, we obtain the below result; however, as was the case for Proposition 15, this could be seen partly as a consequence of our definition of finite for weakly represented families. What the situation looks like in the absence of $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ is left as an open question. We begin with a lemma. It is essentially about purely first-order models, but we phrase it in terms of topped models.
Lemma 20.
Suppose $\mathcal{M}\models\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ is topped by $A$. Let $\Phi^{A}$ be universal in the sense of Lemma 2, suppose $c\in M$ is large enough that $\{e<c:\Phi_{e}^{A}$ is not $\mathcal{M}$-finite}}\}}, and define a partial function $u(e,x)$ by:
$u(e,x)=\min\{s\colon(\forall y<x)\Phi^{A}_{e,s}(y){\downarrow}\}.$
[…]
Proof.
Suppose for a contradiction there is a $\Phi_{e}^{A}$ which is total and such that
$(\forall i<c)(\exists d)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)]$
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the first two quantifiers, this becomes:
$(\exists d)(\forall i<c)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }u(i,x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }u(i,x)<u(e,x)]$
But this means, with knowledge of $d$ and $e$, we the set $\{i<c\colon\Phi_{e}^{A}$ is total}}\}} is $\Pi^{0}_{1}$ and hence $\mathcal{M}$-finite, a contradiction. ∎
Proposition 21.
Every topped model of $\mathsf{RCA}_{0}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}+\neg\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ satisfies $\mathsf{MED}$.
Proof.
Fix $\mathcal{M}$, $A$, $\Phi$, $u$ as in the Lemma. Assume (perturbing the definition if necessary) that $u$ is injective, so that we can list all its values in strictly ascending order:
$u(e_{0},x_{0})<u(e_{1},x_{1})<u(e_{2},x_{2})<\cdots.$
We define by stages a set $(g_{i})_{i<c}$ of partial functions; at stage $s=u(e_{i},x_{i})$, we define $g_{i}(x_{i})$. Begin by forming the set $E=\{e<c\colon\Phi^{A}_{e,s}{\downarrow}$ and $(\forall j\neq i)g_{j}(x)\neq\Phi^{A}_{e,s}(x)\}$. If $E$ is empty, set $g_{i}(x)$ to equal some number that has not been seen previously in the construction. Otherwise, select an $e$ which minimizes $e+|\{y<x\colon\Phi_{e,s}(y)=g_{i}(y)\}|$. This completes the constrution.
We make four claims about $(g_{i})_{i<c}$. The first is that if $i\neq j$, then $g_{i}(x)\neq g_{j}(x)$ for all $x$ which are both defined; this is direct from the construction. The second is that those $g_{i}$ which are total are eventually different; this claim follows a fortiori from the first. The third is that $g_{i}$ is total iff the domain of $g_{i}$ is unbounded in $M$ iff $\Phi_{e}^{A}$ is total; this is direct from the construction and the definition of $u(i,x)$. The fourth claim is that these $g_{i}$ form a maximal such family. For a contradiction, suppose the fourth claim is false, which since the model is topped means that there is some $\Phi^{A}_{e}$ which is total and eventually different from each total $g_{i}$. In symbols:
$(\forall i<c)(\exists d)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }g_{i}(x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }g_{i}(x)\neq\Phi_{e}^{A}(x)].$
Using $\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ to exchange the leading quantifiers, this becomes:
$(\exists d)(\forall i<c)(\forall x>d)[\textnormal{either\leavevmode\nobreak\ }g_{i}(x){\uparrow}\textnormal{\leavevmode\nobreak\ or\leavevmode\nobreak\ }g_{i}(x)\neq\Phi_{e}^{A}(x)].$
Keeping $e$ fixed, let $i$ be as provided by the Lemma. Consider the number $n(x)=e+|\{y<x\colon\Phi^{A}_{e,s}(y)=g_{i}(y)\}|$ which was used in the construction of $g_{i}$. Clearly $n$ is nondecreasing; the equation above tells us is that $n(x)$ reaches some $\mathcal{M}$-finite limit $N$, after some stage $s=\langle i,x\rangle$ of the construction. Then at each stage $t=\langle i,x^{\prime}\rangle$ which is greater than $s$, we know that $\Phi_{e}^{A}(x^{\prime})$ was not chosen to equal $g_{i}(x^{\prime})$. Since $i$ is given by the Lemma, we know that $u(e,x)<u(i,x)$ for $\mathcal{M}$-infinitely many $x$; hence at $\mathcal{M}$-infinitely many stages, there was another $e^{\prime}\in E$ with
$e^{\prime}+|\{y<x\colon\Phi_{e^{\prime},s}(y)=g_{i}(y)\}|<N$
But this is possible only if $e^{\prime}<N$, and then no more than $N-e^{\prime}$ times. This contradicts the pigeonhole principle, and proves the fourth claim. Therefore $(g_{i})_{i<c}$ is a MED family, as required. ∎
7. Nondominated families of functions
The following principle is analogous to the existences of hyperimmune degree.
Statement 22 (Nondominated Family principle, $\mathsf{NDF}$).
Given any weakly represented family $\mathcal{F}$ of functions, there exists a function $g$ such that for all $f\in\mathcal{F}$, there are unboundedly many $x\in M$ such that $g(x)>f(x)$; that is, no $f\in\mathcal{F}$ dominates $g$.
The following statement is immediate.
Proposition 23.
$\mathsf{RCA}_{0}\vdash\mathsf{DOM}\rightarrow\mathsf{NDF}$
More interesting is the question whether the converse holds; the following theorem gives a negative answer. This is achieved by separating $\mathsf{DOM}$ from $\mathsf{NDF}$ by showing that, whereas $\mathsf{DOM}+\mathsf{B}\mathrm{\Sigma}_{2}^{0}$ implies full arithmetical induction as observed in Theorem 6 above, $\mathsf{NDF}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ does not even imply $\mathsf{I}\mathrm{\Sigma}_{3}^{0}$.
Theorem 24.
$\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{NDF}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
Proof.
Fix a model $\mathcal{M}=(M,\mathcal{S})$ of $\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}$ topped by a set $A_{0}\in\mathcal{S}$. Applying the Low Basis Theorem to an $M$-infinite $A_{0}$-computable binary tree without $A_{0}$-computable paths we obtain a set $A_{1}\in\Delta_{2}^{A_{0}}\setminus\Delta_{1}^{A_{0}}$ that is low relative to ${A_{0}}$ and such that $A_{0}\in\Delta_{1}^{A_{1}}$.
Given any $A_{0}$-computable approximation $(A_{1,s})_{s\in M}$ to $A_{1}$, its settling-time $f$, given by
$f(x)=\min\{s\colon(\forall s^{\prime}>s)[f(s^{\prime})=f(s)]\}$
for all $x$, is not dominated by any total $\Delta_{1}^{A_{0}}$ function (as otherwise $A_{1}\in\Delta_{1}^{A_{0}}$ after all). Furthermore, since $A_{1}$ is low relative to ${A_{0}}$, all $\Sigma_{2}^{A_{1}}$ sets are $\Sigma_{1}^{A_{0}^{\prime}}$ and hence regular by Lemma 7. Therefore, $\mathcal{M}_{1}=\mathcal{M}_{0}[A_{1}]$ is a model of $\mathsf{I}\mathrm{\Sigma}_{2}^{0}$.
Repeating the argument, we can obtain $M_{n}$ for all $n<\omega$, and finally we can let $M_{\omega}$ be the model obtained in the limit. Then $\mathcal{M}_{\omega}\models\mathsf{RCA}_{0}+\mathsf{I}\mathrm{\Sigma}_{2}^{0}+\mathsf{NDF}$, as desired. ∎
8. Bi-immunity
Statement 25 (Bi-immunity principle, $\mathsf{BI}$).
For every weakly represented family $\mathcal{F}$ of infinite sets there is a set $B$ such that there is no $A\in\mathcal{F}$ with $A\subseteq B$ or $A\subseteq\overline{B}$.
We can rephrase this statement in terms of the universal sets of Lemma 2 as follows: $\mathsf{BI}$ states that, for every $C$, there exists a $B$ for which there is no infinite $\Delta^{C}_{1}$ set $A=\Phi_{e}^{C}$ such that either $A\subseteq B$ or $A\subseteq\overline{B}$.
The following is a formalization of the well-known fact that every hyperimmune set computes a bi-immune set. Note that, in combination with our previous results, this creates a connection between $\mathsf{BI}$ and $\mathsf{DOM}$, $\mathsf{MAD}$, and induction strength, among others.
Proposition 26.
$\mathsf{RCA}_{0}+\mathsf{NDF}\vdash\mathsf{BI}$.
Proof.
Let $\mathcal{M}=(M,\mathcal{S})$ be a model of $\mathsf{NDF}$, and fix any $A\in\mathcal{S}$. It is sufficient to construct a set $B$ such that no $C=\Phi^{A}_{e}$ is contained in $B$ or $\overline{B}$.
Consider the family of use functions of the functionals $\Phi^{A}_{e}$ for every $e$, and let $f$ witness $\mathsf{NDF}$ with respect to that family. In order to decide which elements we want to put into $B$, we proceed as follows for each $y$ in ascending order: Check if there exists an $e<y$ such that $\Phi^{A}_{e,f(y)}(y){\downarrow}$ but such that for all $x<y$, either $\Phi^{A}_{e,f(y)}(x){\uparrow}$ or $\Phi^{A}_{e,f(y)}(x){\downarrow}=B(x)$. If no such $e$’s exist, then let $y\not\in B$; and if they do exist, then pick the least such $e$ and put $y$ into $B$ or $\overline{B}$ in such way as to ensure $B(y)\neq\Phi^{A}_{e}(y)$. By $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$, every total $\Phi_{e}^{A}$ eventually gets treated in this way. ∎
Finally, here is a simple conservation result.
Theorem 27.
$\mathsf{BI}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}$.
The proof is an imitation of that used by Harrington [6, IX.2] in proving that $\mathsf{WKL}_{0}$ is $\Pi^{1}_{1}$-conservative over $\mathsf{RCA}_{0}$.
Proof.
Suppose that $\mathcal{M}=(M,\mathcal{S})$ is countable model of $\mathsf{RCA}_{0}$ topped by $A\in\mathcal{S}$. If we can construct an $A_{1}\subseteq M$ such that $M_{1}=M[A_{1}]\models\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ and such that $A_{1}$ is bi-immune relative to $A$ (in the sense that for no $e$ is $\Phi_{e}^{A}$ the characteristic function of an $\mathcal{M}$-infinite set which is either a subset of $A_{1}$, or a subset of $\overline{A_{1}}$), then we can iterate the construction to obtain $\mathcal{M}_{2}=\mathcal{M}_{1}[A_{2}]$, and so on, until we obtain in the limit a model $M_{\omega}$ of $\mathsf{RCA}_{0}+\mathsf{BI}$.
We construct $A_{1}$ as the limit of a sequence of initial segments $(\sigma_{i})_{i}$, as follows. To begin, let $(B_{i})_{i<\omega}$ be an enumeration of all $\mathcal{M}$-infinite $A$-computable sets, and let $(c_{i})_{i<\omega}$ be a cofinal sequence in $M$, both ordered by the true natural numbers $\omega$.
  • At stage 00, let $\sigma_{0}$ be the empty string.
  • At stages of the form $s+1=2i+1$ for some $i$, let $\sigma_{s+1}$ be the shortest extension of $\sigma_{s}$ such that $\sigma_{s+1}$ disagrees with (the characteristic function of) $B_{i}$.
  • At stages of the form $s+1=2i+2$ for some $i$, let $\sigma_{s+1}$ be the shortest extension of $\sigma_{s}$ which maximizes the size of $\{e<c_{i}\colon\Phi_{e}(e){\downarrow}\}$.
The odd stages ensure bi-immunity, and the even stages guarantee that $\mathsf{I}\mathrm{\Sigma}_{1}^{0}$ holds. ∎